The buyer for the women’s sweater department has purchased wool sweaters for $35.69. She uses an odd pricing policy and wants to sell them at a 49-percent markup on selling price. At what price should each sweater be sold?
Answer #1

Ans $ 69.98

LET THE SELLING PRICE BE X

COST + MARKUP ON SELLING PRICE = SELLING PRICE

35.69 + 49% OF X = X

35.69 + 0.49X = X

X - 0.49X = 35.69

0.51X = 35.69

X = 35.69 / 0.51

X = $ 69.98